OpenStudy (anonymous):

I have only 2 attempts to get this correct... This is involving a basketball player's three point shooting. Assume that his shots are Bernoulli trials, and that he makes 1 in 7 of his shots. How many three point shots must the player attempt so that the probability of his hitting at least one of them is at least 0.93?

OpenStudy (anonymous):

at least one, means not none. none is \[(\frac{6}{7})^n\] for n trials. so we need \[(\frac{6}{7})^n<.07\]

OpenStudy (anonymous):

i get n = 18 by trial and error

OpenStudy (anonymous):

this will assure that \[1-(\frac{6}{7})^n>.93\] i think we can solve exactly using logarithms, but since n is an integer trial and error seems best

OpenStudy (anonymous):

So try 18

OpenStudy (anonymous):

you can change 18 to a different number and experiment http://www.wolframalpha.com/input/?i=%286%2F7%29^18

OpenStudy (anonymous):

I did some by trial and error but mostly through formulas

OpenStudy (anonymous):

you were right

OpenStudy (anonymous):

imagine!